# Broadband Low-Resolution Spectrograph - SpectrumMate LR

### v2

@@ 2.1 Requirements
In contrast to the original SpectrumMate, which has a spectral coverage of $368\ \textup{\AA}$ in one image (Hien et al.,, 2024), SpectrumMate LR is engineered to capture a broader range of wavelengths, from $3400\ \textup{\AA}$ to $7000\ \textup{\AA}$. This enhanced wavelength range enables the observation of a wide spectrum of celestial features, from the hydrogen and helium lines typical in stellar spectra to more complex molecular signatures found in nebulae and other diffuse objects. The increased bandwidth and optimised resolution make the SpectrumMate LR suitable for capturing both broad spectral features and essential details of emission and absorption lines.
@@
Thirdly, high light sensitivity. The device’s sensitivity must be optimised to perform well with small-aperture telescopes, typically 127 to 254 $mm$ (5 to 10 $inches$) in diameter and focal ratios from about f/5 to f/10, which are common in amateur astronomy. This ensures that SpectrumMate LR can capture sufficient light from fainter objects, such as stars of magnitude 10 or 12, without prohibitively long exposure times, while maintaining a manageable field of view for the spectrograph’s setup.
@@ 2.2 Theoretical optical designs
The theoretical background provided in the work of Buil, (2003, 2024) and Hien et al., (2024) offers a basis for the optical system of SpectrumMate LR (Figure 1), determining optimal spectral range, resolution, and other design parameters to enhance performance for small-telescope compatibility. The two mentioned works also cover telescope aperture compatibility and sensitivity of a spectrograph, which are not covered in this paper.
@@
Distance from the slit to the collimator lens $80\ mm$
Distance from the objective lens to the camera $80\ mm$
Grating groove density $m$ $300\ \textnormal{grooves}/mm$
@@
Slit width $w$ $15\ \mu m$
[…]
@@
Pixel size $p$ $2.4\times 2.4\ \mu m$
@@ 2.2.3 Angles of incidence and diffraction
relates the wavelength of the light $\lambda$, the distance between the grooves $d$, the angle of incidence $\alpha$, the angle of diffraction $\beta$ (both in the unit of $degree$), and the order of the diffraction $k$. It is sometimes convenient to write the equation as
@@
The equations presented are applicable only when the incident and diffracted rays are situated in a plane that is perpendicular to the grooves at the centre of the grating, as shown in Figure 2. This condition, called classical or in-plane diffraction, encompasses most grating systems (Palmer,, 2020).
Figure 2: Grating geometry by Palmer, (2020).
@@
We have $G=\alpha-\beta$, replace $\beta$ with $G$, the grating equation can be written as
@@ 2.2.4 Minimum grating dimension
Since $L<25\ mm$, the chosen holographic diffraction grating is large enough for no light loss to occur.
@@ 2.2.5 Spectral dispersion on sensor
Let $\rho$ be the degree of dispersion in Å$\textup{\AA}/pixel$ on the camera sensor. To optimise data storage, we use $2\times 2$ binning, which groups four neighbouring pixels into one "superpixel" that has the size doubled as the original pixel. Therefore, the effective pixel size $p$ is doubled in dispersion calculations.
@@ 2.2.6 Diameter $d_{2}$ of the diffracted rays
where $X$ is the distance from the grating to the objective lens, and $N$ is the number of pixels across the sensor in the dispersion direction. Due to mechanical constraints, $X$ is set to $80\,mm$. For the chosen sensor (Sony IMX183CLK-J), $N=5496$ pixels.
@@ 2.3 Mechanical Design
[…]
@@ 3.2 Star Classification Based on Spectral Type
We obtained the spectra using the CDK600 telescope at QNO. The mount’s accurate tracking system allowed us to take spectra of even low-brightness objects. This telescope has supported the first detection of exoplanet transit in Vietnam, by Nguyen-Van et al., (2023), and was the main instrument used in SAGI Summer School 2023, detail is shown in the paper written by Nguyen-Luong et al., (2023).
@@ 3.3 Temperature Measurement of Light Sources
The result shows that there is a significant discrepancy between the temperature measured using the peak wavelength and the temperature reported by the manufacturer. This difference can be explained by several factors. First, the peak emission wavelengths observed in the visible range do not correspond to the true blackbody peak for these incandescent light bulbs; the actual emission peaks for the stated temperatures ($2500\ K$ and $3350\ K$) lie in the infrared. Since the spectrograph primarily captures the visible tail of the spectrum, applying Wien’s law to this portion results in an overestimation of the temperature. Second, the calculation assumes the source behaves as an ideal blackbody (with emissivity equal to 1), but real materials, such as tungsten filaments, exhibit wavelength-dependent emissivity that can alter the spectral shape and shift the apparent peak. Finally, instrumental factors such as calibration inaccuracies in the spectrograph’s wavelength scale and sensitivity further contribute to the discrepancy.
@@ 3.4 Filter Transmission Testing
Bandpass $(nm)$ 401-550 562-695 695-844 $>$ 820
[…]
Figure 12: Transmission profile of the filters by the manufacturer. The vertical axis is for transmission (%), and the horizontal axis is for wavelength ($nm$).

### v3

@@ 2.1 Requirements
In contrast to the original SpectrumMate, which has a spectral coverage of $368\ \textup{\AA }$ in one image (phan2024), SpectrumMate LR is engineered to capture a broader range of wavelengths, from $3400\ \textup{\AA }$ to $7000\ \textup{\AA }$. This enhanced wavelength range enables the observation of a wide spectrum of celestial features, from the hydrogen and helium lines typical in stellar spectra to more complex molecular signatures found in nebulae and other diffuse objects. The increased bandwidth and optimised resolution make the SpectrumMate LR suitable for capturing both broad spectral features and essential details of emission and absorption lines.
@@
Thirdly, high light sensitivity. The device’s sensitivity must be optimised to perform well with small-aperture telescopes, typically 127 to 254 $mm$ (5 to 10 inchesinches) in diameter and focal ratios from about f/5 to f/10, which are common in amateur astronomy. This ensures that SpectrumMate LR can capture sufficient light from fainter objects, such as stars of magnitude 10 or 12, without prohibitively long exposure times, while maintaining a manageable field of view for the spectrograph’s setup.
@@ 2.2 Theoretical optical designs
The theoretical background provided in the work of builpara; buil2024 and phan2024 offers a basis for the optical system of SpectrumMate LR (Figure 1), determining optimal spectral range, resolution, and other design parameters to enhance performance for small-telescope compatibility. The two mentioned works also cover telescope aperture compatibility and sensitivity of a spectrograph, which are not covered in this paper.
@@
Distance from the slit to the collimator lens 80mm80\ $m$
Distance from the objective lens to the camera 80mm80\ $m$
Grating groove density $m$ * $300\ \textnormal{grooves}/mm$
@@
Slit width $w$ * $15\ \mu m$
Telescope’s principal mirror diameter $D$ and focal length $f$ * 610mm; 3962mm610\ mm;\ 3962\ $m$
@@
Pixel size $p$ * $2.4\times 2.4\ \mu m$
@@ 2.2.3 Angles of incidence and diffraction
relates the wavelength of the light $\lambda$, the distance between the grooves $d$, the angle of incidence $\alpha$, the angle of diffraction $\beta$ (both in the unit of degreedegree), and the order of the diffraction $k$. It is sometimes convenient to write the equation as
@@
The equations presented are applicable only when the incident and diffracted rays are situated in a plane that is perpendicular to the grooves at the centre of the grating, as shown in Figure 2. This condition, called classical or in-plane diffraction, encompasses most grating systems (palmerdiff).
Figure 2: Grating geometry by palmerdiff.
@@
We have $G=\alpha-\beta\$, replace $\beta$ with $G$, the grating equation can be written as
@@ 2.2.4 Minimum grating dimension
Since L<25mmL<25\ $m$, the chosen holographic diffraction grating is large enough for no light loss to occur.
@@ 2.2.5 Spectral dispersion on sensor
Let $\rho$ be the degree of dispersion in $\textup{\AA }/pixel$ on the camera sensor. To optimise data storage, we use $2\times 2$ binning, which groups four neighbouring pixels into one "superpixel" that has the size doubled as the original pixel. Therefore, the effective pixel size $p$ is doubled in dispersion calculations.
@@ 2.2.6 Diameter $d_{2}$ of the diffracted rays
where $X$ is the distance from the grating to the objective lens, and $N$ is the number of pixels across the sensor in the dispersion direction. Due to mechanical constraints, $X$ is set to 80mm80\,$m$. For the chosen sensor (Sony IMX183CLK-J), $N=5496$ pixels.
@@ 2.3 Mechanical Design
For the collimator lens and its interface, when the lens and slit were installed, the distance between them was maintained at approximately $50\ \textnormal{mm}$, ensuring a robust setup. The objective tube was tested with the ASI183 Pro camera and a focuser, directed at a distant object to simulate parallel rays, successfully achieving focus at infinity. These validations confirm that SpectrumMate LR functions as intended.
@@ 3.2 Star Classification Based on Spectral Type
We obtained the spectra using the CDK600 telescope at QNO. The mount’s accurate tracking system allowed us to take spectra of even low-brightness objects. This telescope has supported the first detection of exoplanet transit in Vietnam, by Tue-Nguyen-Van2023, and was the main instrument used in SAGI Summer School 2023, detail is shown in the paper written by Quang-NGUYEN-LUONG2023.
@@ 3.3 Temperature Measurement of Light Sources
The result shows that there is a significant discrepancy between the temperature measured using the peak wavelength and the temperature reported by the manufacturer. This difference can be explained by several factors. First, the peak emission wavelengths observed in the visible range do not correspond to the true blackbody peak for these incandescent light bulbs; the actual emission peaks for the stated temperatures (2500K2500\ K and 3350K3350\ K) lie in the infrared. Since the spectrograph primarily captures the visible tail of the spectrum, applying Wien’s law to this portion results in an overestimation of the temperature. Second, the calculation assumes the source behaves as an ideal blackbody (with emissivity equal to 1), but real materials, such as tungsten filaments, exhibit wavelength-dependent emissivity that can alter the spectral shape and shift the apparent peak. Finally, instrumental factors such as calibration inaccuracies in the spectrograph’s wavelength scale and sensitivity further contribute to the discrepancy.
@@ 3.4 Filter Transmission Testing
Bandpass $(nm)$ 401-550 562-695 695-844 >> 820
Peak Transmission $(\%)$ 95 98 98 98
Figure 12: Transmission profile of the filters by the manufacturer. The vertical axis is for transmission (%), and the horizontal axis is for wavelength (nmnm).